The density of a given mass of 500g and volume 1000mm³ is given by?
1
Expert's answer
2021-02-19T18:58:05-0500

By definition, the dinsity of the body is given as the ratio of its mass and volume:


"\rho = \dfrac{m}{V}"

Since "m = 500g" is the mass, and "V = 1000mm^3 = 1cm^3" the density will be:


"\rho = \dfrac{500g}{1cm^3} = 500\dfrac{g}{cm^3} = 5\times 10^5\dfrac{kg}{m^3}"

Answer. 500 g/cm^3 or 5*10^5 kg/m^3.